OpenStudy (anonymous):

If six 6-sided fair dice are rolled, what is the probability that three of the dice show prime numbers, and the other three show composite numbers.

OpenStudy (anonymous):

(dice are considered distinguishable)

OpenStudy (anonymous):

hmmm the chance is \[6\] *\[\left(\begin{matrix}6 \\ 3\end{matrix}\right)\]\[\times 0.5^{3}\times 0.5^{3}\]

OpenStudy (anonymous):

Can you please explain your work?

OpenStudy (amistre64):

how many prime numbers on a die?

OpenStudy (anonymous):

There are \[6^{4}\] Which equals to 1296. different results for rolling 4 six-sided dice. The only way the product of the four upper faces can be prime is if three of the upper faces show a 1 and the remaining upper face shows a prime number less than 6, namely 2, 3, or 5. Here are the possibilities 1, 1, 1, 2 1, 1, 1, 3 1, 1, 1, 5 1, 1, 2, 1 1, 1, 3, 1 1, 1, 5, 1 1, 2, 1, 1 1, 3, 1, 1 1, 5, 1, 1 2, 1, 1, 1 3, 1, 1, 1 5, 1, 1, 1 for 12 possible successful outcomes. \[\frac{ 12 }{ 1296 } = \frac{ 1 }{ 108 }\]

OpenStudy (anonymous):

Uhh, I'm not sure that is the answer to the question, it doesn't involve the product of 4 dice.

OpenStudy (anonymous):

Also, I believe there are 4 prime numbers on each die, 1,2,3,5

OpenStudy (amistre64):

1 is not a prime ,,, so lets say 3 out of 6

OpenStudy (anonymous):

Whoops, forgot that 1 only has 1 factor

OpenStudy (amistre64):

lets go with 6 choose 3, 1/2^3 in favor and 1/2^3 not in favor

OpenStudy (anonymous):

Where do I go from there though?

OpenStudy (amistre64):

id multiply ....

OpenStudy (anonymous):

I'm not quite sure what to do.

OpenStudy (amistre64):

6C3 = 20 20 times, 1/2^3 times, 1/2^3 = 20/2^6

OpenStudy (anonymous):

Oh, that's it?

OpenStudy (amistre64):

unless im reading the problem incorrectly ,,, then yes

OpenStudy (anonymous):

I got... 5/16

OpenStudy (amistre64):

me too :)

OpenStudy (anonymous):

It says it's incorrect?

OpenStudy (anonymous):

Its under the topic, "Binomial Probability", if that's any kind of hint.

OpenStudy (amistre64):

thats what we did ...

OpenStudy (amistre64):

6 dice, choose 3 successes 6.5.4 ----- = 20 outcomes for 3primes and 3 composites 3.2.1

OpenStudy (amistre64):

yes it is lol ... 3 primes and 3not hmm, might be overlooking the something; the number of ways a prime can appear

OpenStudy (amistre64):

222 is different from 532

OpenStudy (anonymous):

just to butt in for a second one is NOT prime

OpenStudy (amistre64):

lol, we already established that ;)

OpenStudy (anonymous):

oh did you establish also that this is a simple binomial problem?

OpenStudy (anonymous):

yeah, looks like you did that too back to judge judy for me

OpenStudy (amistre64):

theres like 27runs we can make, but they include repeats 222 223 225 232 233 235 252 253 255 322 323 325 332 333 335 352 353 355 522 523 525 532 533 535 552 553 555 weeding out the repeats 222 223 225 233 235 255 333 335 355 555 looks to be it: 10 of them?

OpenStudy (anonymous):

Then wouldn't you have to multiply by the number of combinations of the composite numbers?

OpenStudy (amistre64):

maybe, ive been staring at this screen way to long today ... brains getting fried :) try the first response; and see if 6*(our result) works

OpenStudy (anonymous):

Dunno, I only have one try left until it is counted wrong :P

OpenStudy (amistre64):

you can afford to miss one right?

OpenStudy (anonymous):

Eh, sorta I guess, the way the system works is, you have to answer a certain amount, but if you miss one, you have to answer I think it's something like... 3-7 more?

OpenStudy (anonymous):

Also, I don't get what you mean by taking about the repeats, can't they stay?

OpenStudy (amistre64):

maybe they can, im trying to consider a smaller "table"

OpenStudy (amistre64):

rolling 2 dice gets(1p, 1c) gets us 1/2 rolling 4 dice tho ... i wonder if its still gonna be 1/2

OpenStudy (mathmate):

There is a problem with the question. 1 is neither a prime nor a composite.

OpenStudy (anonymous):

Yeah, I just now noticed the same thing.

OpenStudy (amistre64):

lol, accursed sets

OpenStudy (anonymous):

So just try what you first did without 1? xD

OpenStudy (amistre64):

for a binomial to work we would have to have p+q = 1 3*3*3*2*2*2 = 216 ways to get 3 primes and 3 composites how many permutation can that create? there are 6*6*6*6*6*6 outcomes total: 46656

OpenStudy (anonymous):

I'm not very good at applying permutations, a lot better at combinations, what would it be, (written like xPy)

OpenStudy (amistre64):

i believe the rule would be:\[\frac{6!}{3!3!}\]

OpenStudy (amistre64):

which is 20 :)

OpenStudy (anonymous):

20

OpenStudy (amistre64):

20(216)/6^6 = 5/54 ... its woth a shot

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

sweet jesus

OpenStudy (anonymous):

Its correct

OpenStudy (amistre64):

lol .... yay!!!!!

OpenStudy (anonymous):

After an hour xD

OpenStudy (anonymous):

Let me check how they did it

OpenStudy (anonymous):

They just did 6c3*(1/2)^3*(1/3)^3

OpenStudy (amistre64):

thats what i thought; but im pretty sure that p+q = 1 in a binomial setup 1/2 + 1/3 is not 1

OpenStudy (amistre64):

maybe a 6, 5 sided dice :) 6C3 (3/5)^3 (2/5)^3 ; but thats off too

OpenStudy (anonymous):

Thanks